# HPVM-HDC: A Heterogeneous Programming System for Accelerating Hyperdimensional Computing

### v2

@@ 1. Introduction
As machine learning models are hitting the memory and compute limits of modern hardware, they are suffering from low performance and high energy consumption on edge devices. In order to meet the performance, energy, and accuracy demands of modern workloads, a brain-inspired computing paradigm, Hyperdimensional Computing (HDC) (Kanerva, 2009), is gaining traction. It represents data using hypervectors, which are high dimensional vectors with thousands of elements. HDC operations are light-weight, highly parallelizable, and robust to noise, so they are well-suited for computations in resource-constrained environments. As a result, HDC has been used in a plethora of applications including natural language processing (Rahimi et al., 2016b), robotics (Mitrokhin et al., 2019), voice and gesture recognition (Imani et al., 2017; Rahimi et al., 2016a), emotion recognition (Chang et al., 2019), graph learning (Nunes et al., 2022), DNA sequencing (Kim et al., 2020), recommender systems (Guo et al., 2021), and bio-signal processing (Asgarinejad et al., 2020).
@@
HDC algorithms can run on multi-core CPUs (Morris et al., 2019; Vergés et al., 2023) and GPUs (Kang et al., 2022b; Simon et al., 2022; Kang et al., 2022a) due to their highly parallel character. In recent years, FPGA-based, ASIC-based, and processing-in-memory (PIM) based HDC accelerators have achieved higher performance and energy efficiency (Yu et al., 2023; Datta et al., 2019; Imani et al., 2019b; Gupta et al., 2018; Imani et al., 2021, 2019d; Salamat et al., 2020; Khaleghi et al., 2021; Yang et al., 2024; Dutta et al., 2022; Xu et al., 2023a; Fan et al., 2024b; Pinge et al., 2024; Zhang et al., 2024; Fan et al., 2024a). These accelerators have specialized on-chip buffers and pipelined components to maximize parallelism and minimize data movement. Some based on resistive RAM (ReRAM) (Xu et al., 2023a) and phase change memory (PCM) (Fan et al., 2024b) also require hardware-specific optimizations to manage trade-offs between accuracy and energy efficiency / performance.
@@
Custom accelerators will only succeed if they are supported by practical, widely adopted programming models and development tools. In the case of HDC, realistic applications must be able to run on custom accelerators as well as more widely available devices like CPUs and GPUs because suitable accelerators are usually available on some systems and not others. Even when an accelerator is available, in our experience, accelerators only support a proper subset of core HDC operations, forcing some application components to run on other hardware; moreover, this partitioning between accelerator and general purpose hardware is different for different accelerators. To achieve this kind of portability and flexible application partitioning, HDC applications require appropriate high-level abstractions to effectively target accelerators and heterogeneous hardware. Without such abstractions, application developers must write their applications to run well on each specific target because an application parallelized for GPU cannot be recompiled and run for an HDC ASIC or for a PIM-based device. Writing such multi-version code would be impractical for production software.
@@
There are at least three challenges in designing and supporting high-level, easily portable programming abstractions for HDC applications. First, HDC programming requires novel and unusual data encoding schemes, combined with hypervector and hypermatrix operations that depend on those encoding schemes (Heddes et al., 2022). These custom schemes must be captured in the high-level programming model without putting undue burden on the programmer to implement the encoding-specific details of common operations. Second, to achieve high performance on many targets, including CPUs, GPUs, and others, programs must use hardware-specific optimizations that maximize data locality and leverage different forms of parallelism in HDC algorithms. High-level abstractions cannot expose these kinds of optimization options or hardware primitives. Instead, compilers must automatically achieve such optimization requirements while minimizing the need for hardware-specific tuning and programmer guidance (e.g., using command line options or configuration files). Third, achieving adequate accuracy with HDC algorithms is difficult and hardware-dependent: it requires optimizing different algorithmic choices, including the encoding strategy (data-dependent encoding, random projection, custom implementation, etc.), the similarity search schemes (cosine similarity, Hamming distance, etc.), the number of iterations, hypervector length and element precision, and different approximation choices. The optimal choices often vary between hardware devices, forcing difficult hardware-specific tuning. Again, compilers must automatically perform such tuning with limited programmer guidance.
@@
Prior work in HDC algorithm development use general purpose and target specific languages such as Python (Van Rossum et al., 1995; Tuomanen, 2018; Simon et al., 2022; Heddes et al., 2022; Kang et al., 2022b) and MATLAB (Matlab, 2012; Fatica and Jeong, 2007), but rely on CUDA for GPUs (Sanders and Kandrot, 2010; Kim et al., 2020), and C++ with hardware-specific primitives for FPGAs (Salamat et al., 2019; Imani et al., 2019d). HDCC (Vergés et al., 2023) is a HDC-specific compiler that can compile high-level descriptions of HDC algorithms to multi-threaded C code with vector extensions, but it only targets multi-core CPUs. Applications targeting each of the devices previously mentioned are often split into entirely separate programs, possibly in different languages. Programmers must implement and hand-optimize different versions of HDC algorithms using different languages to target various devices, significantly hampering programmer productivity. Moreover, targeting HDC accelerators that support specialized coarse-grained instructions requires writing target-specific code that is very different from the lower-level operations used on CPUs and GPUs.
@@
To address the challenges described above, we propose a C++-based programming language, HDC++ (Section 3), that provides three major benefits. First, it creates a set of common high-level abstractions specifically for HDC that enable programmers to implement HDC algorithms without worrying about low-level target-specific optimizations. Second, it strikes a balance between the complexity of implementing HDC algorithms and flexibility in exploring different algorithmic and approximation choices by providing high-level abstractions that make tweaking algorithms easier while also allowing programmers to implement custom algorithms for their HDC applications. (Section 5.4). Third, it improves programmer productivity by abstracting away the hardware executing the application—the provided primitives are hardware-independent, so the same HDC++ application can be compiled and executed on CPUs, GPUs, and multiple different HDC accelerators without code modification. We have ported five applications to HDC++ and successfully compiled them to suitable hardware targets. We show in Section 5.2 that using a hardware-agnostic language does not have a negative impact on performance.
@@
We also propose HPVM-HDC (Section 4), a compiler based on Heterogeneous Parallel Virtual Machine (HPVM) (Kotsifakou et al., 2018; Ejjeh et al., 2022) that compiles hardware agnostic HDC++ applications automatically to several hardware targets, including CPUs, GPUs, and two custom HDC accelerators. HDC++ programs are lowered to a retargetable intermediate representation (IR) with HDC-specific operations. The HPVM-HDC IR structurally captures the parallelism in HDC algorithms, enabling compilation onto hardware targets with coarse-grained instructions such as HDC accelerators as well as targets exposing fine-grained parallelism such as GPUs. Additionally, HPVM-HDC IR supports domain-specific optimizations that exploit the inherent trade-off in HDC between end-to-end quality of service and performance. We implement two such optimizations: automatic binarization, which enables programmers to automatically modify hypervector lengths and precisions to use efficient 1-bit level computations, and reduction perforation, which skips elements in hypervectors when computing reductions. Such approximations lead to significant speedups with little programmer effort and marginal loss in application-level quality of service (Section 5.3). HPVM-HDC lowers HDC-specific operations through target specific backends for a wide variety of hardware targets: CPUs, GPUs, and two custom hardware accelerators: a taped-out digital ASIC and a simulated Resistive RAM (ReRAM) accelerator (Section 4.3). Baseline implementations of our evaluated HDC applications exist only for CPUs and GPUs (and are target-specific)—HPVM-HDC achieves competitive performance with those baseline implementations using a single retargetable implementation, achieving a geomean speed-up of 1.17x against optimized baseline CUDA implementations on GPUs. For the accelerators we compile applications to using HPVM-HDC, no prior evaluation has been performed with a full HDC application.
@@
  • The first retargetable compilation framework for HDC, HPVM-HDC, which compiles HDC++ programs and generates high-performance code for a variety of hardware targets including CPUs, GPUs, an ASIC-based accelerator, and a ReRAM-based accelerator simulator (Section 4).
@@ 2.1. Hyperdimensional Computing
Hyperdimensional Computing (HDC) is a machine learning paradigm inspired by cognitive models of computation. HDC algorithms operate by encoding input features into a high dimensional representation called hypervectors. In contrast to Deep Neural Networks (DNNs) which rely on expensive feed-forward computations for inference, HDC relies on lightweight search through the class hypervector database representing the model. The most similar class hypervector signifies the predicted label. HDC training uses efficient element-wise addition instead of back-propagation. Fine-tuning HDC hyper-parameters (such as element representation, encoding dimensionality and training iterations) and algorithm choices play a much more important role in practice for achieving good quality with HDC in comparison to DNNs, where back-propagation can compensate for sub-optimal choices during training. HDC algorithms generally vary in the used encoding schemes, search algorithms, and training procedures which are often application specific. These different algorithms can be used to implement many kinds of applications, including classification (Zhang et al., 2023), clustering (Imani et al., 2019a), and GNN learning (Kang et al., 2022d) (see Table 2). The lightweight and parallel nature of HDC enables high performance across hardware targets.
@@
Figure 1. Architecture block diagrams for the HD Digital ASIC (left) and HD ReRAM accelerator (right).
@@ 2.2. HDC Hardware Accelerators
HPVM-HDC compiles to two custom HDC accelerators. In this section we briefly describe the hardware architectures. Figure 1 shows diagrams for these architectures.
HD Digital ASIC: We run HDC workloads on an ASIC chip fabricated in a 40nm technology. The architecture of the ASIC is described in (Yang et al., 2024) and uses pipelined encoding and Hamming distance search. The ASIC provides a functional interface for device configuration, data movement, and training and inference operations. Input data points are encoded automatically when moved to the device. See Listing 6 for example code generated by HPVM-HDC targeting this interface. Execution times are measured on the chip.
HD ReRAM: We used a simulator for a resistive RAM (ReRAM) accelerator to emulate the performance of HDC code executing on ReRAM hardware. The design is described in (Xu et al., 2023a), and has a similar programming interface to the digital ASIC chip. The simulator uses extracted timing and energy parameters from commercial SRAM and ReRAM macros in the 40nm technology node provided by a foundry. These parameters are loaded into the HD ReRAM simulator (executed on a host machine) to estimate the runtime.
@@ 2.3. HPVM Compiler Infrastructure
The HPVM Compiler Infrastructure (Ejjeh et al., 2022) extends the LLVM compiler infrastructure with a retargetable parallel program representation for heterogeneous architectures (as shown in Figure 2). HPVM IR provides constructs for representing various forms of parallelism including task-level, data-level, and streaming parallelism. (shown in Figure 3). Parallel programs are represented using a hierarchical directed acyclic graph (DAG). Nodes in the graph are categorized into two categories: internal nodes, which capture hierarchical parallelism by containing entire sub-graphs, and leaf nodes, which represent individual units of computation. Edges between nodes in the HPVM DAG represent logical data transfers. Additionally, each node is annotated with one or more hardware targets, and code is generated by the appropriate back ends for each of those targets.
@@ 2.4. Hetero-C++
Hetero-C++ is a parallel C++ dialect which compiles to HPVM IR (Ejjeh et al., 2022). Hetero-C++ provides a high-level, target-agnostic programming interface for capturing the various forms of parallelism in their application. A function in Hetero-C++ may be split into multiple parallel tasks and parallel loops which individually would be lowered to distinct HPVM IR nodes. These distinct components in a single application are individually compiled to different targets as specified by the user, easing programmability for applications on heterogeneous systems and improving portability across systems. Both HDC++ and Hetero-C++ use marker functions instead of pragmas to retain placement within the emitted LLVM IR.
@@ 3. The HDC++ Language
HDC++ is a domain specific language built on top of Hetero-C++, a parallel dialect of C++ (Ejjeh et al., 2022). (Note that the choice of this base language is not essential: any widely used compiled language with explicit task and data parallelism would achieve the same purpose.) HDC++ introduces primitives for creating and manipulating hypervectors and hypermatrices, allowing developers to easily develop HDC algorithms inside their applications. Listing 1 shows how HDC++ can be used to express classification inference. First, random projection encoding is used to encode input features into a 2048-dimensional representation. Then, Hamming distance is applied between the encoded hypervector and the class hypervectors (represented using a hypermatrix datatype) to produce a vector of dissimilarities. In this example, the result is a 26 dimensional vector, as there are 26 output label classes. Finally, the HDC arg_min operation identifies the corresponding class hypervector which is most similar to the encoded hypervector.
Listing 1: HDC++ example for random-projection encoding and dis-similarity between query and class hypervectors.
In addition to HDC specific primitives, application developers writing HDC++ can use Hetero-C++ primitives for expressing generic parallel code. HDC applications exhibit parallelism both within the HDC algorithmic primitives (i.e. operator data-level parallelism), as well as across operations (i.e. task-level or outer-loop-level parallelism). By enabling HDC++ to interoperate with Hetero-C++, application developers can compactly describe complex HDC algorithms using HDC primitives while sufficiently describing any other parallelism using Hetero-C++ and keeping the application entirely retargetable. For example, the HyperOMS (Kang et al., 2022c) application uses level ID encoding, which has an outer parallel loop that isn’t captured by HDC primitives, but is captured by Hetero-C++’s parallel loop annotation. HDC++ can express all the components of the HyperOMS application in a parallel fashion, providing both programmability and performance.

### v3

@@ Abstract.
Compilers, Hyperdimensional Computing, Heterogeneous Systems
journalyear: 2025copyright: cccopyright: ccconference: Proceedings of the 52nd Annual International Symposium on Computer Architecture; June 21–25, 2025; Tokyo, Japanbooktitle: Proceedings of the 52nd Annual International Symposium on Computer Architecture (ISCA ’25), June 21–25, 2025, Tokyo, Japandoi: 10.1145/3695053.3731095isbn: 979-8-4007-1261-6/2025/06ccs: Computer systems organization Heterogeneous (hybrid) systemsccs: Software and its engineering Domain specific languages
@@ 1. Introduction
As machine learning models are hitting the memory and compute limits of modern hardware, they are suffering from low performance and high energy consumption on edge devices. In order to meet the performance, energy, and accuracy demands of modern workloads, a brain-inspired computing paradigm, Hyperdimensional Computing (HDC) (Kanerva, 2009), is gaining traction. It represents data using hypervectors, which are high dimensional vectors with thousands of elements. HDC operations are light-weight, highly parallelizable, and robust to noise, so they are well-suited for computations in resource-constrained environments. As a result, HDC has been used in a plethora of applications including natural language processing (Rahimi et al., 2016b), robotics (Mitrokhin et al., 2019), voice and gesture recognition (Imani et al., 2017; Rahimi et al., 2016a), emotion recognition (Chang et al., 2019), graph and hypergraph learning (Nunes et al., 2022; Kang et al., 2024), DNA sequencing (Kim et al., 2020), recommender systems (Guo et al., 2021), and bio-signal processing (Asgarinejad et al., 2020).
@@
HDC algorithms can run on multi-core CPUs (Morris et al., 2019; Vergés et al., 2023) and GPUs (Kang et al., 2022b; Simon et al., 2022; Kang et al., 2022a) due to their highly parallel character. In recent years, FPGA-based, ASIC-based, and processing-in-memory (PIM) based HDC accelerators have achieved higher performance and energy efficiency (Yu et al., 2023; Datta et al., 2019; Imani et al., 2019b; Gupta et al., 2018; Imani et al., 2021, 2019d; Salamat et al., 2020; Khaleghi et al., 2021; Yang et al., 2024; Dutta et al., 2022; Xu et al., 2023a; Fan et al., 2024b; Pinge et al., 2024; Zhang et al., 2024; Fan et al., 2024a). These accelerators have specialized on-chip buffers and pipelined components to maximize parallelism and minimize data movement. Some devices based on resistive RAM (ReRAM) (Xu et al., 2023a) and phase change memory (PCM) (Fan et al., 2024b) also require hardware-specific optimizations to manage trade-offs between accuracy and energy efficiency / performance.
@@
Custom accelerators will only succeed if they are supported by practical, widely adopted programming models and development tools. In the case of HDC, realistic applications must be able to run on custom accelerators as well as more widely available devices, like CPUs and GPUs, because suitable accelerators are not universally available. Even when an accelerator is available, in our experience, accelerators only support a proper subset of core HDC operations, forcing some application components to run on other hardware; moreover, this partitioning between accelerator and general purpose hardware is different for different accelerators. To achieve this kind of portability and flexible application partitioning, HDC applications require appropriate high-level abstractions to effectively target heterogeneous systems. Without such abstractions, application developers must write their applications to run well on each specific target because an application parallelized for GPU cannot be recompiled and run for an HDC ASIC or for a PIM-based device. Writing such multi-version code would be impractical for production software.
@@
There are at least three challenges in designing and supporting high-level, easily portable programming abstractions for HDC applications. First, HDC programming requires novel and unusual data encoding schemes, combined with hypervector and hypermatrix operations that depend on those encoding schemes (Heddes et al., 2022). These custom schemes must be captured in the high-level programming model without putting undue burden on the programmer to implement the encoding-specific details of common operations. Second, to achieve high performance on many targets, including CPUs, GPUs, and others, programs must use hardware-specific optimizations that maximize data locality and leverage different forms of parallelism in HDC algorithms. High-level abstractions cannot expose these kinds of optimization options or hardware primitives. Instead, compilers must automatically apply these optimizations with minimal programmer guidance. Third, achieving adequate accuracy with HDC algorithms is difficult and hardware-dependent: it requires exploring different algorithmic choices, including the encoding strategy (data-dependent encoding, random projection, custom implementation, etc.), the similarity search schemes (cosine similarity, Hamming distance, etc.), the number of iterations, hypervector length and element precision, and different approximation choices. The optimal choices may vary or be restricted on certain HDC accelerators, forcing difficult hardware-specific tuning. Again, compilers must automatically perform such tuning with limited programmer guidance.
@@
Prior works in HDC algorithm development use general purpose languages such as Python (Van Rossum et al., 1995; Tuomanen, 2018; Simon et al., 2022; Heddes et al., 2022; Kang et al., 2022b) and MATLAB (Matlab, 2012; Fatica and Jeong, 2007), and rely on target specific languages for programming accelerators, such as CUDA for GPUs (Sanders and Kandrot, 2010; Kim et al., 2020), and C++ with hardware-specific primitives for FPGAs (Salamat et al., 2019; Imani et al., 2019d). HDCC (Vergés et al., 2023) is a HDC-specific compiler that can compile high-level descriptions of HDC algorithms to multi-threaded C code with vector extensions, but it’s programming model is not expressive and it only targets multi-core CPUs. Applications targeting each of the devices previously mentioned are often split into entirely separate programs, possibly in different languages. Programmers must implement and hand-optimize different versions of HDC algorithms using different languages to target various devices, significantly hampering programmer productivity. Moreover, targeting HDC accelerators that support specialized coarse-grained instructions requires writing target-specific code that is very different from the lower-level operations used on CPUs and GPUs.
@@
To address the challenges described above, we propose a C++-based programming language, HDC++ (Section 3), that provides three major benefits. First, it creates a set of common high-level abstractions specifically for HDC that enable programmers to implement HDC algorithms without worrying about low-level target-specific optimizations. Second, it strikes a balance between the complexity of implementing HDC algorithms and flexibility in exploring different algorithmic and approximation choices by providing high-level abstractions that make tweaking algorithms easier while also allowing programmers to implement custom algorithms for their HDC applications. Third, it improves programmer productivity by abstracting away the hardware executing the application—the provided primitives are hardware-independent, so the same HDC++ application can be compiled and executed on CPUs, GPUs, and multiple different HDC accelerators without code modification. We have ported five applications to HDC++.
@@
We also propose HPVM-HDC (Section 4), a compiler based on Heterogeneous Parallel Virtual Machine (HPVM) (Kotsifakou et al., 2018; Ejjeh et al., 2022) that compiles hardware agnostic HDC++ applications automatically onto CPUs, GPUs, a digital HDC ASIC and a simulated ReRAM HDC accelerator. HDC++ programs are lowered to a retargetable intermediate representation (IR) with HDC-specific operations (Section 4.1). The HPVM-HDC IR structurally captures the parallelism in HDC algorithms, enabling compilation onto hardware targets with coarse-grained instructions such as HDC accelerators as well as targets exposing fine-grained parallelism such as GPUs. Additionally, HPVM-HDC IR supports domain-specific optimizations that exploit the inherent trade-off in HDC between end-to-end quality of service and performance. We implement two such optimizations: automatic binarization, which enables programmers to automatically modify hypervector lengths and precisions to use efficient 1-bit level computations, and reduction perforation, which skips elements in hypervectors when computing reductions. Such approximations lead to significant speedups with little programmer effort and marginal loss in application-level quality of service (Section 4.2). HPVM-HDC lowers HDC-specific operations through target specific backends for a wide variety of hardware targets: CPUs, GPUs, a taped-out digital HDC ASIC, and a simulated Resistive RAM (ReRAM) HDC accelerator (Section 4.3). Baseline implementations of our evaluated HDC applications exist only for CPUs and GPUs (and are target-specific)—HPVM-HDC achieves competitive performance with those baseline implementations using a single retargetable implementation, achieving a geomean speed-up of 1.17x against optimized baseline CUDA implementations on GPUs. For the accelerators we compile applications to using HPVM-HDC, no prior evaluation has been performed with a full HDC application.
@@
  • The first retargetable compilation framework for HDC, HPVM-HDC, which compiles HDC++ programs and generates high-performance code for a variety of hardware targets including CPUs, GPUs, a digital HDC ASIC, and a simulated ReRAM HDC accelerator (Section 4).
@@ 2.1. Hyperdimensional Computing
Hyperdimensional Computing (HDC) is a machine learning paradigm inspired by cognitive models of computation. HDC algorithms operate by encoding input features into a high dimensional representation called hypervectors. In contrast to Deep Neural Networks (DNNs) which rely on expensive feed-forward computations for inference, HDC relies on encoding query vectors into a high dimensional space and searching through a class hypervector database representing the model—the most similar class hypervector signifies the predicted label. Both of these operations are extremely lightweight in comparison to repeated matrix multiplications. Additionally, HDC training uses repeated element-wise addition instead of back-propagation.
For several tasks, such as graph and hypergraph learning (Nunes et al., 2022; Kang et al., 2024) and federated learning (Tian et al., 2025), HDC techniques have been proposed that achieve significant improvements in latency and energy efficiency over prior DNN-based approaches while maintaining competitive quality of service. While these improvements are significant when deployed on GPU-based systems, HDC solutions have been shown to achieve orders of magnitude improvements in power usage when deployed on specialized hardware (Kang et al., 2024; Yang et al., 2024; Xu et al., 2023a). Additionally, prior work combines HDC techniques with DNN approaches to achieve both high accuracy and energy efficiency (Xu et al., 2023a; Tian et al., 2025).
Fine-tuning HDC hyper-parameters (such as element representation, encoding dimensionality and training iterations) and algorithm choices play a much more important role in practice for achieving good quality with HDC in comparison to DNNs, where back-propagation can compensate for sub-optimal choices during training. HDC algorithms generally vary in the used encoding schemes, search algorithms, and training procedures which are often application specific. These different algorithms can be used to implement many kinds of applications, including classification (Zhang et al., 2023), clustering (Imani et al., 2019a), and GNN learning (Kang et al., 2022d) (see Table 2). The lightweight and parallel nature of HDC enables high performance across hardware targets.
@@
Figure 1. Architecture block diagrams for the digital HDC ASIC (left) and ReRAM HDC accelerator (right).
@@ 2.2. HDC Hardware Accelerators
HPVM-HDC compiles to two custom HDC accelerators. In this section we briefly describe the devices. Figure 1 shows diagrams for their architectures.
HD Digital ASIC: We run HDC workloads on an ASIC chip fabricated in a 40nm technology. The architecture of the ASIC is described by Yang et al. (Yang et al., 2024). The device supports cyclic random projecting operations for encoding and pipelined Hamming distance for both training and inference. In this work, we just use the HDC module, which achieves 0.78 TOPS/W (6.6x higher than prior state-of-the-art). The ASIC provides a functional interface for device configuration, data movement, training, and inference. An FPGA is directly wired to the ASIC and an ARM CPU to facilitate communication—the ARM CPU acts as the host and the ASIC has its own private memory. See Listing 6 for example code generated by HPVM-HDC targeting this interface. Execution times are measured on the chip.
HD ReRAM: We also run HDC programs on a resistive RAM (ReRAM) based HDC accelerator. The architecture of the accelerator is described by Xu et al. (Xu et al., 2023a), and has a similar programming interface to the digital ASIC. The device supports ”tensorized” encoding, which is a more energy efficient variant of random projection encoding, summation based one-shot training, and Hamming distance for inference. Hamming distances are progressively computed until the relative ranking between candidate hypervectors can no longer change. The encoding and inference stages are accelerated using a large ReRAM array. We used a simulator of the device to emulate the performance of HDC code executing on ReRAM hardware. The simulator uses extracted timing and energy parameters from commercial SRAM and ReRAM macros in the 40nm technology node provided by a foundry.
@@ 2.3. HPVM Compiler Infrastructure
The HPVM Compiler Infrastructure (Ejjeh et al., 2022) extends the LLVM compiler infrastructure with a retargetable parallel program representation for heterogeneous architectures (as shown in Figure 2). HPVM IR provides constructs for representing various forms of parallelism including task-level, data-level, and streaming parallelism (and expresses vector parallelism in the LLVM code). Parallel programs are represented using a hierarchical directed acyclic graph (DAG). Nodes in the graph fall into two categories: internal nodes, which capture hierarchical parallelism by containing entire sub-graphs, and leaf nodes, which represent individual units of computation, represented as LLVM IR (shown in Figure 3). Edges between nodes in the HPVM DAG represent logical data transfers, i.e., explicit data copies may or may not be needed. Additionally, each node is annotated with one or more hardware targets, and code is generated by the appropriate back ends for each of those targets (possibly multiple targets).
@@ 2.4. Hetero-C++
Hetero-C++ is a parallel C++ dialect, with parallel features similar to a subset of OpenMP (primarily for ease of porting C++ applications without a full OpenMP front-end). Hetero-C++ compiles to HPVM IR (Ejjeh et al., 2022). Hetero-C++ provides a high-level, target-agnostic programming interface for capturing the various forms of parallelism in an application. A function in Hetero-C++ may be split into multiple parallel tasks and parallel loops which are individually lowered to distinct HPVM IR nodes. These distinct components in a single application are individually compiled to different targets as specified by the user, easing programmability for applications on heterogeneous systems and improving portability across systems. Both HDC++ and Hetero-C++ use marker functions instead of pragmas to retain placement within the emitted LLVM IR.
@@ 3. The HDC++ Language
HDC++ is a domain specific language built on top of Hetero-C++ (Ejjeh et al., 2022). Note that the choice of this base language is not essential in our work: any compiled parallel language with explicit task and data parallelism would achieve the same purpose. HDC++ introduces primitives for creating and manipulating hypervectors and hypermatrices, allowing developers to easily develop HDC algorithms inside their applications. Listing 1 shows how HDC++ can be used to express classification inference. First, random projection encoding is used to encode input features into a 2048-dimensional representation. Then, Hamming distance is applied between the encoded hypervector and the class hypervectors (represented using a hypermatrix datatype) to produce a vector of dissimilarities. In this example, the result is a 26 dimensional vector, as there are 26 output label classes. Finally, the HDC arg_min operation identifies the corresponding class hypervector which is least dissimilar to the encoded hypervector.
1hypervector<617> input_features = ...;
2hypermatrix<2048, 617> rp_matrix = ...;
3hypermatrix<26, 2048> clusters_matrix = ...;
4hypervector<2048> encoded_data = __hetero_hdc_matmul(input_features, rp_matrix);
5hypervector<26> dists = __hetero_hdc_hamming_distance(encoded_data, clusters_matrix);
6int label = __hetero_hdc_arg_min(dists);
Listing 1: HDC++ example for random-projection encoding and dissimilarity between query and class hypervectors.
In addition to HDC specific primitives, application developers writing HDC++ can use Hetero-C++ primitives for expressing generic parallel code. HDC applications exhibit parallelism both within the HDC algorithmic primitives (operator data-level parallelism), as well as across operations (task-level or outer-loop-level parallelism). By enabling HDC++ to interoperate with Hetero-C++, application developers can compactly describe complex HDC algorithms using HDC primitives while describing inter-operator parallelism using Hetero-C++ and keeping the application retargetable. For example, the HyperOMS (Kang et al., 2022c) application uses level ID encoding, which has an outer parallel loop that isn’t captured by HDC primitives, but is captured by Hetero-C++’s parallel loop annotation. HDC++ can express all of the components of the HyperOMS application in a parallel fashion, providing both programmability and performance.